# Human Data-Driven Calibration of Hand Exoskeletons Using Redundant Sensors for Improved Teleoperation

### v1

@@ Abstract
Hand exoskeletons are critical tools for dexterous teleoperation and immersive manipulation interfaces, but achieving accurate hand tracking remains a challenge due to user-specific anatomical variability and donning inconsistencies. These issues lead to kinematic misalignments that degrade tracking performance and limit applicability in precision tasks. We propose a subject-specific calibration framework for exoskeleton-based hand tracking that uses redundant joint sensing and a residual-weighted optimization strategy to estimate virtual link parameters. Implemented on the Maestro exoskeleton, our method improves joint angle and fingertip position estimation across users with varying hand geometries. We introduce a data-driven approach to empirically tune cost function weights using motion capture ground truth, enabling more accurate and consistent calibration across participants. Quantitative results from seven subjects show substantial reductions in joint and fingertip tracking errors compared to uncalibrated and evenly weighted models. Qualitative visualizations using a Unity-based virtual hand further confirm improvements in motion fidelity. The proposed framework generalizes across exoskeleton designs with closed-loop kinematics and minimal sensing, and lays the foundation for high-fidelity teleoperation and learning-from-demonstration applications.
Index Terms:
Hand Tracking, Hand Exoskeleton, Dexterous Manipulation, Kinematics.
@@ I Introduction
Dexterous teleoperation, where humans remotely control robotic hands, offers a promising pathway toward this goal. It is essential for applications such as robotic surgery, space exploration, and disaster response, as well as for collecting high-quality demonstrations to train robots via imitation learning [4]. In both real-time control and demonstration settings, the effectiveness of teleoperation hinges on accurately capturing the user’s hand movements, driving growing interest in evaluating and improving hand tracking methods for these applications.
Current approaches fall into two main categories: vision-based systems and wearable devices [5, 6]. Vision-based methods, including optical motion capture and camera-based frameworks like MediaPipe [7], offer non-invasive, real-time tracking and can achieve sub-centimeter fingertip precision under ideal conditions. However, they remain sensitive to occlusions, lighting variability, and camera placement, which compromises tracking stability and usable workspace [8]. Additionally, they lack haptic feedback, making them less suitable for force-sensitive teleoperation tasks [9].
In contrast, wearable devices—–such as sensorized gloves and hand exoskeletons–—offer a more direct means of capturing finger motion through joint angle or fingertip sensing [10]. Among these, hand exoskeletons are particularly effective in cluttered or occlusion-prone environments [6, 11], and can deliver haptic feedback to keep users actively engaged during dexterous teleoperation [12]. However, exoskeletons face a persistent challenge in maintaining anatomically accurate tracking across diverse users and donning conditions [13]. Variability in hand size, donning inconsistencies, and slippage at the skin–device interface often introduce kinematic misalignments that degrade tracking accuracy and cause discomfort or unintended interaction forces [14].
To mitigate these issues, prior work has employed redundant passive mechanisms [15, 16, 17] or underactuated fingertip-only designs [18, 19]. While these designs enable self-alignment to accommodate donning variability, they do not guarantee accurate joint-level tracking. The introduced kinematic redundancy breaks the one-to-one mapping between exoskeleton and anatomical joints, making the system sensitive to slippage, donning variability, and hand size. As a result, even with high-quality sensors, the estimated joint angles can diverge from the user’s true finger posture, making precise hand tracking difficult [20].
To address this gap, we propose a subject-specific calibration framework aimed at reducing hand tracking error in exoskeleton-based telemanipulation. Our method estimates virtual link parameters that account for different hand sizes and donning conditions, thereby improving the accuracy of hand tracking across users.
@@
Specifically, we implement and validate this approach on the Maestro hand exoskeleton , shown inFig. 1, which employs closed-loop four-bar linkages for passive self-alignment. While this design improves comfort and wearability, it also introduces kinematic redundancy, making the system highly sensitive to hand size, hand anatomy, slippage, and donning conditions [14]. Our proposed calibration mitigates these effects by optimizing virtual link lengths and weighting factors to align exoskeleton outputs with anatomical motion. To address these challenges, we introduce a subject-specific, data-driven calibration framework that aligns exoskeleton outputs with anatomical motion by optimizing virtual link lengths and residual weightings. Instead of relying on rigid hardware alignment, the method estimates these parameters from user motion data, compensating for variations in hand size, finger proportions, and donning configuration.
@@
The calibration is formulated as an optimization problem that minimizes tracking error via a residual-weighted cost function combining joint and fingertip errors. We further introduce a data-driven approach to optimize these weights, enabling the calibration to adapt to subject-specific biomechanics and sensor contributions. This integrated strategy leads to more accurate hand posture estimation and improved tracking fidelity. We implement and evaluate the framework on the Maestro hand exoskeleton, demonstrating consistent improvements in joint and fingertip tracking across seven participants with diverse hand geometries. The method is validated both quantitatively and qualitatively, using a virtual hand model rendered in Unity to assess visual plausibility in teleoperation.
Our key contributions are: (i) a subject-specific calibration procedure for virtual link estimation under anatomical and donning variability; (ii) a residual-weight optimization method that enhances tracking accuracy over traditional uniform weighting; and (iii) a comprehensive multi-subject evaluation and virtual validation demonstrating improved fidelity for teleoperation. Together, these contributions establish a generalizable calibration framework for improving exoskeleton-based hand tracking. While demonstrated on Maestro, the method applies broadly to other wearable hand exoskeletons with kinematic misalignment or minimal sensing, supporting high-fidelity control in teleoperation and learning-from-demonstration applications.
@@ II-A Hardware Platform
(a) Thumb Kinematic Model
(b) Index (and Middle) Finger Kinematic Model
Figure 2: Kinematic diagram showing the four-bar mechanisms implemented in the human-robot kinematic model of the Maestro hand exoskeleton. (a) The thumb uses three kinematic loops, whereas (b) the index and middle fingers use two kinematic loops. The first loop is an RRPR mechanism, and both the second and third loops are RRRR mechanisms. Specific definitions of each variable are provided in Tab. I.
The Maestro hand exoskeleton, shown in Fig.1, is used in this study as a representative case to present and validate our subject-specific calibration framework. Maestro enables motion tracking and haptic feedback for the thumb, index, and middle fingers via a cable-driven[21, 22]. It includes 8 active and 8 passive joints, each instrumented with rotary potentiometers for joint sensing. The index and middle fingers are each equipped with 5 sensors (2 providing redundant measurements), while the thumb has 6 sensors, including two redundant ones. As shown in the figure, the system can measure or estimate the MCP and PIP joints for the fingers, and the CMC, MCP, and IP joints for the thumb. The distal interphalangeal (DIP) joints are computed from the PIP joints using established biomechanical correlations.
@@ II-B Kinematic Model
In the following, we present the human–exoskeleton kinematic model, which maps the exoskeleton’s sensorized joint angles to estimated human joint angles. For the thumb, the sensorized exoskeleton joint angles are denoted as $(\alpha_{2},\beta_{2},\gamma_{2},\delta_{1},\delta_{2},\delta_{3})$, and the corresponding anatomical joint angles are $(\theta_{1},\theta_{2},\theta_{3})$, representing the CMC, MCP, and IP joints, respectively. While similar models apply to the thumb, index, and middle fingers, here we only detail the thumb model for clarity. Complete loop equations and joint mappings are provided in Fig. 2 and Tab. I.
@@
As mentioned above, the exoskeleton includes redundant sensor inputs, which can also be derived from the geometry of the kinematic loops. These redundant joint angles are incorporated into the calibration cost function to improve estimation accuracy. Here, $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ represent redundant joint angles derived from specific four-bar link constraints, as shown in Fig. 2.
Each redundant joint encodes constraints across one or more loops: i) $\hat{\delta}_{1}$ is defined as the angle between links $b_{1}$ and $d_{2}$ and depends on both the first and second loops; ii) $\hat{\delta}_{2}$, defined as the angle between $a_{2}$ and the projection of $b_{2}$ along the $y$-axis ($y_{3}$), is specific to the index and middle fingers and is influenced by the second loop geometry, $f(\beta)$; and iii) $\hat{\delta}_{3}$ the angle between $a_{2}$ and $d_{3}$ is unique to the thumb, and depends on both the second and third loops, $f(\beta,\gamma)$.
@@
$\begin{split}\hat{\delta}_{3}=2\pi-\beta_{4}-(\pi/2-\beta_{6})-\gamma_{3}-(\pi/2-\gamma_{1})\end{split}$ (3)
where the intermediate variables are presented in Tab. I.
@@
II-C Kinematic Parameters Sensitivity Analysis
To motivate the need for a calibration procedure that adjusts the kinematic parameters values, we first analyzed the sensitivity of fingertip position to perturbations in the virtual link parameters. Using the kinematics model of the index finger, each of the six 2D virtual link coordinates $\{x_{1},y_{1},x_{2},y_{2},x_{3},y_{3}\}$ was perturbed independently in simulation, and the resulting deviation in fingertip position was computed. The goal was to demonstrate that even small errors in these parameters can lead to significant inaccuracies in fingertip tracking, thereby justifying the need for precise calibration.
The results, shown in Fig. 3, indicate that proximal parameters—particularly $x_{1}$ and $x_{3}$—have a significantly larger effect on fingertip position than more distal ones such as $y_{1}$ and $y_{3}$. This confirms that small errors in estimating virtual link parameters can lead to substantial errors in fingertip position—up to 30 $\mathrm{m}\mathrm{m}$ for just a 10% perturbation in proximal coordinates.
Furthermore, due to the kinematic design, variations in horizontal parameters (i.e., $x$ values) tend to produce larger fingertip tracking errors than equivalent changes in vertical ones ($y$ values). This is particularly relevant given that horizontal sliding of the device on the back of the hand is more likely than vertical shifting during wear. Such horizontal misalignments directly impact the inferred finger pose, emphasizing the need for personalized calibration of virtual link positions to maintain accurate fingertip estimation.
[Parameter $x_{1}$] Refer to caption [Parameter $x_{3}$] Refer to caption
Figure 3: Sensitivity analysis in simulation of fingertip position error (mm) as a function of perturbations in select parameters (variation).
@@ II-C Kinematic Parameters Sensitivity Analysis
Input: $\alpha_{2}$ Input: $\beta_{2},\delta_{1},\delta_{2}$ Input: $\gamma_{2},\delta_{3}$
@@
Output: $\theta_{1}$ Output: $\theta_{2},\hat{\delta}_{1},\hat{\delta}_{2}$ Output: $\theta_{3},\hat{\delta}_{3}$
@@
$\begin{array}[]{rl}d_{1}&=\sqrt{\scriptstyle a_{1}^{2}+b_{1}^{2}-c_{1}^{2}+2a_{1}b_{1}\cos(\alpha_{2}-\alpha_{1})}\\[3.44444pt] \alpha_{3}&=-2\arctan\left(\tiny{\frac{\begin{array}[]{@{}c@{}}b_{1}-2d_{1}+2a_{1}\cos(\alpha_{2})\cos(\alpha_{1})\\ +2(b_{1}-d_{1})\cos(\alpha_{2})+2a_{1}\cos(\alpha_{1})\\ +b_{1}\cos(2\alpha_{2})\end{array}}{\begin{array}[]{@{}c@{}}(2\cos(\alpha_{2})+1)\cdot\\ \left(-c_{1}+a_{1}\sin(\alpha_{1})+b_{1}\sin(\alpha_{2})\scriptsize\right)\end{array}}}\right)\end{array}$ $\begin{array}[]{rl}c_{2}&=\sqrt{\scriptstyle{c_{1}^{2}+(L_{1}-d_{1})^{2}}}\\[3.44444pt] \beta_{1}&=\arctan\left(\frac{c_{1}}{L_{1}-d_{1}}\right)\\[3.44444pt] \beta_{3}&=\arcsin\left(\frac{a_{2}\sin(\beta_{2})}{\sqrt{a_{2}^{2}+d_{2}^{2}-2a_{2}d_{2}\cos(\beta_{2})}}\right)\\[10.33333pt] \beta_{4}&=\arccos\left(\frac{c_{2}^{2}+d_{2}^{2}-(a_{2}^{2}+b_{2}^{2})-2c_{2}d_{2}\cos(\beta_{2})}{-2a_{2}b_{2}}\right)\\[10.33333pt] \beta_{5}&=\arccos\left(\frac{a_{2}^{2}+d_{2}^{2}-(b_{2}^{2}+c_{2}^{2})-2a_{2}d_{2}\cos(\beta_{2})}{-2b_{2}c_{2}}\right)\\[6.88889pt] &+\arcsin\left(\frac{b_{2}\sin(\beta_{6})}{\sqrt{b_{2}^{2}+c_{2}^{2}-2b_{2}c_{2}\cos(\beta_{6})}}\right)\par\end{array}$ $\begin{array}[]{rl}\gamma_{3}&=\arcsin\left(\frac{a_{3}\cdot\sin(\gamma_{2})}{\sqrt{a_{3}^{2}+d_{3}^{2}-2a_{3}d_{3}\cos(\gamma_{2})}}\right)+\\ &+\arcsin\left(\frac{b_{3}\cdot\sin(\gamma_{5})}{\sqrt{b_{3}^{2}+c_{3}^{2}-2b_{3}c_{3}\cos(\gamma_{5})}}\right)\\ \gamma_{5}&=\arccos\left(\frac{a_{3}^{2}+d_{3}^{2}-(b_{3}^{2}+c_{3}^{2})-2a_{3}d_{3}\cdot\cos(\gamma_{2})}{-2b_{3}c_{3}}\right)\\ \end{array}$
@@
II-D Human-Exoskeleton Kinematic Calibration
@@
The goal of the calibration process is to optimize subject- and donning-specific virtual link parameters (e.g., a1, c1, b2, c3) by minimizing discrepancies between estimated ($\hat{\theta}_{n}$, $\hat{\delta}_{n}$) and measured joint angles ($\delta_{n,\text{ref}}$) or known configurations ($\theta_{n,\text{ref}}$). The inputs to the calibration are the measured joint angles, while the outputs are the estimated virtual links, which generate updated estimates of the user’s joint angles.
The full procedure, illustrated in Fig. 4,consists of two phases:
@@
In phase one ($\phi_{1}$), subjects fully extend their fingers to form a flat-hand posture and hold it for a few seconds while static joint measurements are recorded. This configuration imposes known anatomical constraints, which serve as reference angles ($\theta_{\text{ref}}$ and $\delta_{\text{ref}}$) for calibration: all finger joints are assumed to be at 0°0$\degree$0 °, except for the thumb CMC flexion/extension, which is anatomically set at 70°70$\degree$70 °.
In phase two ($\phi_{2}$), subjects perform isolated flexion of the MCP joints of the thumb, index, and middle fingers while keeping the IP (thumb) and PIP (index/middle) joints extended. In this case, the reference angles for the MCP joints are unknown and serve as targets for optimization, while the IP/PIP joints are again assumed to remain at 0°0$\degree$0 °.
@@
There are two interchangeable representations of virtual link parameters: the length–angle form $(a_{1},\alpha_{1},c_{2},\beta_{1},b_{2},\beta_{6},b_{3},\gamma_{6})$, used in the kinematic model, and the XY-coordinate form $(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3},x_{4},y_{4})$, used as input for optimization. The conversion between the two is done via trigonometric identities. We adopt the length–angle form for computing joint angles using closed-form geometric relations, and the XY-coordinate form for optimization due to its dimensional consistency.
The calibration cost function is defined as a weighted sum of squared errors between estimated joint angles and either known reference or redundant sensor readings, respectively for phase one and two. We define each error term $\Delta$ as the discrepancy between the model and ground truth, as follows:
@@
$f_{1}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}\Big{(}$ $w_{1}(\Delta\theta_{1,\phi_{1}}^{(i)})^{2}+w_{2}(\Delta\theta_{2,\phi_{1}}^{(i)})^{2}+w_{3}(\Delta\delta_{1,\phi_{1}}^{(i)})^{2}+$ (9)
@@
where $\mathbf{x}=(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})$, $w_{k}$ are weights and $N$ is the number of samples. To ensure a well-posed optimization, the number of error terms must match the number of virtual link parameters. For example, six parameters require six independent error terms. While both $\Delta\delta_{1,\phi_{2}}$ and $\Delta\delta_{2,\phi_{2}}$ are valid choices, we empirically found that using $\Delta\delta_{1,\phi_{2}}$ results in better calibration.
@@
$f_{2}(\mathbf{y})=\frac{1}{N}\sum_{i=1}^{N}\Big{(}w_{1}(\Delta\theta_{1,\phi_{1}}^{(i)})^{2}+w_{2}(\Delta\theta_{2,\phi_{1}}^{(i)})^{2}+w_{3}(\Delta\theta_{3,\phi_{1}}^{(i)})^{2}+$ (10)
$w_{4}(\Delta\delta_{1,\phi_{1}}^{(i)})^{2}+w_{5}(\Delta\delta_{2,\phi_{1}}^{(i)})^{2}+w_{6}(\Delta\theta_{3,\phi_{2}}^{(i)})^{2}+$
$w_{7}(\Delta\delta_{1,\phi_{2}}^{(i)})^{2}+w_{8}(\Delta\delta_{2,\phi_{2}}^{(i)})^{2}\Big{)}$
where $\mathbf{y}=(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3},x_{4},y_{4})$. Subsequently, the global optima of the two problems are obtained as:
$\mathbf{x}^{*}=\arg\min_{\mathbf{x}}f_{1}(\mathbf{x}),\quad\mathbf{y}^{*}=\arg\min_{\mathbf{y}}f_{2}(\mathbf{y})$ (11)
The kinematic calibration procedure relies on optimizing a set of virtual link parameters by minimizing a weighted sum of joint angle errors. The performance of this optimization depends not only on the quality of the input data but also on how the error terms are weighted in the cost function. While the simulation-based sensitivity analysis, presented in II-C, offers a first-principles rationale for prioritizing proximal parameters, it does not capture real-world factors such as model inaccuracies and inter-subject variability. To address this, we conducted a data-driven weight refinement via motion capture validation (see Section III-A).

### v2

# Human-Exoskeleton Kinematic Calibration to Improve Hand Tracking for Dexterous Teleoperation

@@ Abstract
Hand exoskeletons are critical tools for dexterous teleoperation and immersive manipulation interfaces, but achieving accurate hand tracking remains a challenge due to user-specific anatomical variability and donning inconsistencies. These issues lead to kinematic misalignments that degrade tracking performance and limit applicability in precision tasks. We propose a subject-specific calibration framework for exoskeleton-based hand tracking that estimates virtual link parameters through residual-weighted optimization. A data-driven approach is introduced to empirically tune cost function weights using motion capture ground truth, enabling accurate and consistent calibration across users. Implemented on the Maestro hand exoskeleton with seven healthy participants, the method achieved substantial reductions in joint and fingertip tracking errors across diverse hand geometries. Qualitative visualizations using a Unity-based virtual hand further demonstrate improved motion fidelity. The proposed framework generalizes to exoskeletons with closed-loop kinematics and minimal sensing, laying the foundation for high-fidelity teleoperation and robot learning applications.
@@ I Introduction
Dexterous teleoperation, where humans control robotic hands to perform fine tasks, offers a practical pathway toward this goal, supporting applications from surgery and space exploration to robot learning from human demonstration [4]. Its success depends critically on accurate, low-latency hand tracking that reliably captures user motion.
Existing approaches fall into two main categories: vision-based systems and wearable devices [5, 6]. Vision-based methods such as optical motion capture or camera-based tracking [7, 8], provide non-invasive sensing but suffer from occlusions, lighting sensitivity, and workspace limits, and they lack haptic feedback for force-based interactions [9].
@@
In contrast, wearable devices, such as data gloves or exoskeletons, directly measure joint angles or fingertip motion [10, 11, 12], enabling operation in cluttered environments and providing force feedback. However, their accuracy declines due to user-specific anatomy, inconsistent donning, and slippage [13, 14] and resulting in discomfort in mechanical designs that directly align device joints with anatomical joints [15].
Self-alignment mechanisms [16, 17] mitigate kinematic incompatibility but often introduce non-linear correspondence between device and anatomy. To complement these mechanical solutions, recent research has focused on calibration strategies that explicitly compensate for anatomical variability, donning inconsistencies, and sensor misalignment [18].
Commercial devices still rely on manual, pose-based calibration routines that are labor-intensive, user-dependent, and can take up to 90 minutes per glove, limiting practicality. Heinrich et al. [19] highlighted the trade-off between calibration accuracy and complexity, motivating faster, more generalizable approaches. Recent work by Li et al. [20] further advanced shape- and subject-specific calibration through unified kinematic–anatomical models that enhance tracking accuracy and consistency across users.
Hybrid and sensor-fusion approaches have also been explored. Gosala et al. [21] combined visual and magnetic sensors, reducing joint-angle error to 11.4° but requiring both cameras and wearables. Kim et al. [22] improved fingertip accuracy via data-driven optimization, though it depended on subject-specific ground-truth data.
Building on these insights, our framework estimates subject-specific kinematic parameters to establish an individualized mapping between the hand and the exoskeleton. This mapping is derived from redundant sensing and a limited set of reference poses, without requiring external cameras, sensors or ground-truth references.
@@
Specifically, we implement and validate this approach on the Maestro hand exoskeleton (Fig. 1), which employs self-aligning closed-loop four-bar linkages for kinematic compatibility. While this architecture enhances comfort and wearability, it remains sensitive to differences in hand size, anatomy, slippage, and donning variability [14]. To address these effects, we develop a subject-specific, data-driven calibration framework that estimates virtual link parameters and residual weightings directly from user motion data, compensating for geometric and anatomical variability without requiring rigid mechanical alignment.
@@
The calibration is formulated as an optimization problem that minimizes tracking error through a residual-weighted cost function combining joint and fingertip discrepancies. The data-driven weighting scheme allows adaptation to subject-specific biomechanics and sensor contributions, resulting in more accurate hand-pose estimation and improved tracking fidelity. We evaluate the framework on seven participants with diverse hand geometries and validate its performance both quantitatively and qualitatively using a Unity-based virtual hand visualization. Although demonstrated on the Maestro exoskeleton, the proposed framework establishes a general foundation for accurate, subject-specific calibration in sensorized hand exoskeletons used for teleoperation and learning-from-demonstration.
@@ II-A Hardware Platform
The Maestro hand exoskeleton, shown in Fig. 1, is used in this study as a representative case to present and validate our subject-specific calibration framework. Maestro enables motion tracking and haptic feedback for the thumb, index, and middle fingers via a cable-driven mechanism [23, 24]. The device comprises 16 joints instrumented with rotary potentiometers: five sensors per index and middle finger and six for the thumb, each including two redundant measurements.
@@ II-B Kinematic Model
In the following, we present the human–exoskeleton kinematic model that maps the exoskeleton’s sensorized joint angles to the corresponding anatomical joint angles. For the index and middle fingers, the joint angles are $(\alpha_{2},\beta_{2},\delta_{1},\delta_{2})$, where $\delta_{1}$ and $\delta_{2}$ represent the two redundant joints. Whereas for the thumb, the sensorized exoskeleton joint angles are denoted as $(\alpha_{2},\beta_{2},\gamma_{2},\delta_{1},\delta_{2}^{\prime})$, where $\delta_{1}$ and $\delta_{2}^{\prime}$ serve as redundant measurements.The corresponding anatomical joint angles are $(\theta_{1},\theta_{2})$ for the MCP and PIP joints of the index and middle fingers, and $(\theta_{1},\theta_{2},\theta_{3})$ for the CMC, MCP, and IP joints of the thumb, respectively. Although similar models are defined for all digits, the thumb model is presented here for clarity. Complete loop equations and joint mappings are provided in Fig. 2 and Tab. I.
@@
The exoskeleton includes redundant joint sensors derived from its kinematic loops, incorporated into the calibration cost to improve estimation accuracy. Each digit has two redundant sensors: $(\delta_{1},\delta_{2}^{\prime})$ for index/middle and $(\delta_{1},\delta_{2})$ for the thumb, as shown in Fig. 2.
The first redundant joint, $\hat{\delta}_{1}$, is common to all digits and is defined as the angle between links $b_{1}$ and $d_{2}$, depending on both the first and second loops. The second redundant joint differs by digit: for the index and middle fingers, $\hat{\delta}_{2}$ is defined as the angle between $a_{2}$ and the projection of $b_{2}$ along the $y$-axis ($y_{3}$), influenced by the second loop geometry $f(\beta)$; for the thumb, $\hat{\delta}_{2}^{\prime}$ is defined as the angle between $a_{2}$ and $d_{3}$, depending on the second and third loops $f(\beta,\gamma)$. Thus, $\delta_{2}$ denotes the second redundant joint of the index/middle fingers, while $\delta_{2}^{\prime}$ denotes the second redundant joint of the thumb.
@@
$\begin{split}\hat{\delta}_{2}^{\prime}=2\pi-\beta_{4}-(\pi/2-\beta_{6})-\gamma_{3}-(\pi/2-\gamma_{1})\end{split}$ (3)
where the intermediate variables are presented in Tab. I. These angles, estimated from loop geometry, provide additional constraints used in the calibration cost function.
@@
(a) Thumb Kinematic Model
(b) Index (and Middle) Finger Kinematic Model
Figure 2: Kinematic model of the Maestro finger–exoskeleton four-bar interface. The thumb is equipped with redundant joints $(\delta_{1},\delta_{2}^{\prime})$ and is modeled with three kinematic loops: one RRPR loop and two RRRR loops. The index and middle fingers are equipped with redundant joints $(\delta_{1},\delta_{2})$ and are modeled with two kinematic loops: one RRPR loop and one RRRR loop. The XXYY frame is defined such that the $X$-axis is aligned longitudinally with the base frame and the $Y$-axis is perpendicular to it. All variable definitions are provided in Tab. I.
@@
Input: $\alpha_{2}$ Input: $\beta_{2},\delta_{1},\delta_{2}$ Input: $\gamma_{2},\delta_{2}^{\prime}$
@@
Output: $\theta_{1}$ Output: $\theta_{2},\hat{\delta}_{1},\hat{\delta}_{2}$ (index/middle only) Output: $\theta_{3},\hat{\delta}_{2}^{\prime}$ (thumb only)
@@
[…]
@@
II-C Kinematic Parameters Sensitivity Analysis
To motivate the need for a calibration procedure that adjusts the kinematic parameter values, we first analyzed the sensitivity of fingertip position to perturbations in the virtual link parameters. The goal was to demonstrate that even small errors in these parameters can lead to significant inaccuracies in fingertip tracking. Using the kinematic model of the index finger, each of the six 2D virtual link coordinates $\{x_{1},y_{1},x_{2},y_{2},x_{3},y_{3}\}$ was perturbed independently within $\pm$ 10% of its nominal value in simulation, and the resulting deviation in fingertip position was computed.
The results, shown in Fig. 3, indicate that proximal parameters, particularly $x_{1}$ and $y_{1}$, exert the strongest influence on fingertip position, with 10% perturbations causing deviations of up to 30 $\mathrm{m}\mathrm{m}$. In contrast, distal parameters such as $x_{3}$ and $y_{3}$ had only minor effects, typically below 3–5 $\mathrm{m}\mathrm{m}$.
Fingertip tracking was more sensitive to horizontal ($x$) perturbations, consistent with common horizontal sliding of the device on the back of the hand during use. This finding underscores the need for calibration that optimizes virtual link parameters to maintain fingertip tracking accuracy.
[Parameter $x_{1}$] Refer to caption [Parameter $x_{2}$] Refer to caption [Parameter $x_{3}$] Refer to caption [Parameter $y_{1}$] Refer to caption [Parameter $y_{2}$] Refer to caption [Parameter $y_{3}$] Refer to caption
Figure 3: Sensitivity analysis in simulation of fingertip position error ($m$) as a function of perturbations in select parameters (% variation).
II-D Human-Exoskeleton Calibration Framework
@@
The goal of the calibration process is to optimize subject- and donning-specific virtual link parameters (e.g., a1, c1, b2, c3) by minimizing discrepancies between estimated ($\hat{\theta}_{n}$, $\hat{\delta}_{n}$) and measured joint angles ($\delta_{n,\text{ref}}$) or known configurations ($\theta_{n,\text{ref}}$). The inputs to the calibration are the measured joint angles, while the outputs are the estimated virtual links, which generate updated estimates of the user’s joint angles. The full procedure is illustrated in Fig. 4.
@@
In phase one ($\phi_{1}$), subjects fully extend their fingers to form a flat-hand posture and hold it for a few seconds while static joint measurements are recorded. This configuration imposes known anatomical constraints, which serve as reference angles ($\theta_{\text{ref}}$ and $\delta_{\text{ref}}$) for calibration: all finger joints are assumed to be at $0\degree$, except for the thumb CMC flexion/extension, which we fixed at 70°when participants opened their hands in a flat posture. Although this value is not anatomically identical across individuals, it provides a consistent reference across subjects and ensures the calibration problem remains well-posed.
In phase two ($\phi_{2}$), subjects perform isolated flexion of the MCP joints of the thumb, index, and middle fingers while keeping the IP (thumb) and PIP (index/middle) joints extended. Here, the MCP reference angles are unknown and serve as optimization targets, while the IP/PIP joints are again assumed to remain at $0\degree$.
@@
Virtual link parameters can be expressed in two equivalent forms: a length–angle form $(a_{1},\alpha_{1},c_{2},\beta_{1},b_{2},\beta_{6},b_{3},\gamma_{6})$ for kinematic modeling and an XY-coordinate form $(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3},x_{4},y_{4})$ for optimization. The two are related through trigonometric identities; the former enables closed-form angle computation, while the latter ensures dimensional consistency during optimization.
The calibration minimizes a weighted sum of squared errors between model-predicted and reference joint angles, where each error term $\Delta$ represents the discrepancy between estimated and measured values for anatomical or redundant joints, respectively (Eqs. 7-8).
@@
$f_{1}(\mathbf{p})=\frac{1}{N}\sum_{i=1}^{N}\Big($ $w_{1}(\Delta\theta_{1,\phi_{1}}^{(i)})^{2}+w_{2}(\Delta\theta_{2,\phi_{1}}^{(i)})^{2}+w_{3}(\Delta\delta_{1,\phi_{1}}^{(i)})^{2}+$ (9)
@@
where $\mathbf{p}=(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})$, $w_{k}$ are weights and $N$ is the number of samples. To ensure a well-posed optimization, the number of error terms must match the number of virtual link parameters. This condition ensures that each parameter can be uniquely estimated from independent error constraints without introducing underdetermined or redundant solutions.
To ensure a well-posed optimization, the number of error terms must match the number of virtual link parameters. For example, six parameters require six independent error terms. Although both $\Delta\delta_{1,\phi_{2}}$ and $\Delta\delta_{2,\phi_{2}}$ are theoretically valid, $\Delta\delta_{1,\phi_{2}}$ was used since $\delta_{2,\phi_{2}}$ was prone to interface tilting and inconsistent contact during flexion.
@@
$f_{2}(\mathbf{q})=\frac{1}{N}\sum_{i=1}^{N}\Big(w_{1}(\Delta\theta_{1,\phi_{1}}^{(i)})^{2}+w_{2}(\Delta\theta_{2,\phi_{1}}^{(i)})^{2}+w_{3}(\Delta\theta_{3,\phi_{1}}^{(i)})^{2}+$ (10)
$w_{4}(\Delta\delta_{1,\phi_{1}}^{(i)})^{2}+w_{5}(\Delta\delta_{2,\phi_{1}}^{\prime(i)})^{2}+w_{6}(\Delta\theta_{3,\phi_{2}}^{(i)})^{2}+$
$w_{7}(\Delta\delta_{1,\phi_{2}}^{(i)})^{2}+w_{8}(\Delta\delta_{2,\phi_{2}}^{\prime(i)})^{2}\Big)$
where $\mathbf{q}=(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3},x_{4},y_{4})$. Subsequently, the global optima of the two problems are obtained as:
$\mathbf{p}^{*}=\arg\min_{\mathbf{p}}f_{1}(\mathbf{p}),\quad\mathbf{q}^{*}=\arg\min_{\mathbf{q}}f_{2}(\mathbf{q})$ (11)
The kinematic calibration procedure relies on optimizing a set of virtual link parameters by minimizing a weighted sum of joint angle errors with the solver fminuncfminunc from MATLAB. The performance of this optimization depends not only on the quality of the input data but also on how the error terms are weighted in the cost function. While the simulation-based sensitivity analysis, presented in II-C, offers a first-principles rationale for prioritizing proximal parameters, it does not capture real-world factors such as model inaccuracies and inter-subject variability. To address this, we performed a data-driven weight refinement using motion-capture validation (see Section III-A).
Although illustrated on the Maestro four-bar design, the same calibration principle applies to other hand-exoskeleton architectures, provided that the system includes a sufficient number of independent constraints, whether from redundant robot sensors, geometric closures or constrained hand reference poses, to identify the unknown kinematic parameters. These constraints guarantee that the optimization remains well-posed; further discussion on generalizability is provided in Section V.